# Radiative Corrections to Light Thermal Pseudo-Dirac Dark Matter

### v1

@@ Abstract
Light thermal dark matter has emerged as an attractive theoretical possibility and a promising target for discovery at experiments in the near future. Such scenarios generically invoke mediators with very small couplings to the Standard Model, but moderately strong couplings within the dark sector, calling into question theoretical estimates based on the lowest order of perturbation theory. As an example, we focus on a scenario in which (pseudo)-Dirac fermion dark matter is connected to the standard model via a dark photon charged under a new $U(1)^{\prime}$ extension of the standard model, and we investigate the impact of the next-to-leading order corrections to annihilation and scattering. We find that radiative corrections can significantly impact model predictions for the relic density and scattering cross-section, depending on the strength of the dark sector coupling and ratio of the dark matter to mediator mass. We also show why factorization into the yield parameter $Y$ typically presented in literature leads to imprecision. Our results are necessary to accurately map experimental searches into the model parameter space and assess their ability to reach thermal production targets.
@@ I Introduction
Of all of the viable scenarios of Dark Matter (DM), one of the most compelling is that of particles that are in thermal equilibrium with the Standard Model (SM) bath in the early Universe. DM with a thermal history is very well motivated because it not only provides a feasible prediction for large non-gravitational interactions between dark and ordinary matter, but is also highly predictive in nature, leading to clear targets for experimental searches. Generically, after the DM has frozen out and its relic density set, it maintains the same interaction strength with the SM today, predicting that it is likely to be observable at a variety of experiments on Earth Bertone and Tait (2018). A canonical example is that of Weakly interacting massive particles (WIMPs), a class of heavy ($\sim$ GeV - TeV scale) particles that interact with the SM through a roughly electroweak strength force Cushman et al. (2013); Cooley et al. (2022).
@@
However, despite decades of search, heavy ($\gtrsim$ GeV) WIMPs are yet to be discovered and their parameter space has become tightly constrained Cooley et al. (2022). This lack of a clear signal has motivated searches for other visions of DM, including models with masses in the MeV to GeV regime interacting with the SM via new undiscovered forces. While this class of DM is no longer a WIMP in the traditional sense, it can still be in equilibrium at early times and thus represents a thermal target. It is a very attractive prospect as it may provide a road map to a wide and rich dark sector. The possibilities of simplified DM models are vast. However, one of the most appealing scenarios is that of DM interacting with the SM through a new vector boson, often refered to as a dark photon, corresponding to a $U(1)^{\prime}$ extension of the SM Fayet (1980, 1981); Holdom (1986).
@@
Thermal sub-GeV DM, through a dark photon portal is very well motivated and has been the subject of much exploration in the literature. In this scenario, MeV to GeV mass DM can be produced and detected at a variety of current and near-future low energy accelerator experiments Izaguirre et al. (2015a, b, c, 2016); Feng and Smolinsky (2017); Izaguirre et al. (2017); Aguilar-Arevalo et al. (2018); Marsicano et al. (2018); Berlin and Kling (2019); Mohlabeng (2019); Tsai et al. (2021); Duerr et al. (2020); Berlin et al. (2020); Batell et al. (2021a, b); Carrillo González and Toro (2022); Cazzaniga et al. (2021); Gori et al. (2022); Krnjaic et al. (2022); Mongillo et al. (2023); Abdullahi et al. (2023); Brahma et al. (2023). Moreover, ambient DM in the Solar System can scatter with electrons and nucleons in small scale direct detection experiments with low energy recoil thresholds Emken et al. (2019); Bloch et al. (2021); Essig et al. (2022). In addition, accelerator and underground detector probes are well complemented by astrophysical and cosmological constraints An et al. (2018); Chang et al. (2018); DeRocco et al. (2019); Fitzpatrick et al. (2022); Giovanetti et al. (2022). All together, these probes make for an exciting experimental program that shows promise for discovering light thermal DM in the near future.
@@
As an illustrative model, we focus on pseudo-Dirac dark matter interacting with the SM via a dark photon mediator. Current experimental constraints and projections on this model that target the dark matter thermal relic abundance focus on a dark sector coupling $\alpha_{D}$ that is stronger than the QCD coupling at the electroweak scale. As a result, the dark sector is strongly coupled and relatively large corrections from higher orders of perturbation theory are expected. In this work, we compute the next-to-leading (NLO) order corrections on both the thermal annihilation and late time scattering processes with electrons. We focus on the NLO corrections at ${\cal O}(\alpha_{D})$ and neglect the presumably negligible higher order corrections from the tiny kinetic mixing.
@@
We find that the NLO corrections can be as large as ${\cal O}(10\%)$ for parameters typically discussed in the literature, and are thus necessary to take into account when precisely mapping experimental searches into the parameter space, and when comparing them with thermal production milestones which serve as a prime target for a currently growing intensity frontier program. It is worth pointing out that various theoretical constraints on this model resulting from its strong coupling have been considered in the literature, e.g. from the running of the dark sector coupling and the breakdown of perturbation theory Davoudiasl and Marciano (2015); Reilly and Toro (2023).
@@
The rest of this article is organized as follows: In section II we describe the reference model, section III is devoted to a full description of the thermal annihilation cross-section, including both leading and next-to-leading order (NLO) processes. We compute the scattering cross-sections with electrons relevant for direct detection searches in section IV. Finally we conclude in section V. The appendix provides some technical details related to the counter-terms in the on-shell renormalization scheme.
@@ II Pseudo-Dirac Dark Matter
The basic module we consider consists of two Weyl fermions which play the role of dark matter and are paired by a Dirac mass, $m_{\chi}$. These fermions are neutral under the SM gauge groups, but have equal and opposite charge ±$\pm 1$ under a gauged dark U(1), with corresponding gauge boson $A^{\prime}_{\mu}$. The U(1) symmetry is spontaneously broken by the vacuum expectation value of a dark Higgs scalar $\phi$, generating a mass for $A^{\prime}_{\mu}$ and (assuming the $\phi$ charge under U(1) is chosen appropriately) Majorana masses for the two Weyl fermions. We further assume that there is kinetic mixing between the dark photon and the SM hypercharge interaction, induced by unspecified UV physics.
@@
where $\chi$ is a Dirac fermion packaging both of the original Weyl fermions in the limit of zero mass splitting (and thus including both the dark matter and its heavier partner when the mass splitting is taken into account). $D_{\mu}\equiv\partial_{\mu}-ig_{D}Q_{A^{\prime}}A^{\prime}_{\mu}$ is the covariant derivative for field of U(1) charge $Q_{A^{\prime}}$, $X_{\mu\nu}$ and $B_{\mu\nu}$ are the field strengths for the U(1) and SM hypercharge bosons, respectively, $\varepsilon$ characterizes the strength of the kinetic mixing and $\phi$ is the dark Higgs field. We normalize $Q_{A^{\prime}}=1$ from here on for $\chi$. After symmetry breaking, the scalar field acquires a vacuum expectation value (VEV) $v_{\phi}$ and can be parameterized in the unitary gauge as $\phi\rightarrow(v_{\phi}+H_{D})/\sqrt{2}$, where $H_{D}$ is the dark Higgs boson.
@@ III Thermal Freeze out
For dark matter masses below $\lesssim$ GeV, the annihilation rate into hadrons may be inferred by making use of the fact that $A^{\prime}$ couples to the same electromagnetic current as the photon. Thus, the effective cross section for annihilation into hadrons at a given center-of-mass-energy can be related to the ratio $R$ measured in $e^{+}e^{-}$ reactions Olive et al. (2014):
@@
III.1 Tree level annihilation
At tree level, the annihilation into SM fermions proceeds via $s$-channel exchange of the $A^{\prime}$ (see Figure 1). The leading order matrix element $\overline{|{\cal M}|^{2}}_{\text{LO}}$ is given by,
@@ III.1 Tree level annihilation
$\Gamma(A^{\prime}\rightarrow\chi\bar{\chi})=\frac{1}{3}\,\alpha_{D}\,m_{A^{\prime}}\left(1+\frac{2m_{\chi}^{2}}{m_{A^{\prime}}^{2}}\right)\,\sqrt{1-\frac{4m_{\chi}^{2}}{m_{A^{\prime}}^{2}}},$ (6)
$\Gamma(A^{\prime}\rightarrow f\bar{f})=\frac{1}{3}\,\varepsilon^{2}\alpha\,m_{A^{\prime}}\left(1+\frac{2m_{f}^{2}}{m_{A^{\prime}}^{2}}\right)\,\sqrt{1-\frac{4m_{f}^{2}}{m_{A^{\prime}}^{2}}},$ (7)
@@
In the non-relativistic limit, the thermally-averaged tree-level cross section is
@@ III.2 NLO Corrections
At ${\cal O}(\alpha_{D}^{2})$, the annihilation cross section receives virtual corrections in the form of self-energy corrections to the dark photon propagator and incoming dark matter wave functions and a correction to the $A^{\prime}$-$\chi$-$\overline{\chi}$ vertex (see Figure 2). In the regime of $m_{A}^{\prime}\geq 2m_{\chi}$, an additional emitted $A^{\prime}$ would always be off-shell, leading the real emission corrections to be effectively higher order in $\varepsilon$ such that they can be safely neglected. The ${\cal O}(\alpha_{D}^{2})$ correction to the annihilation process is given by the interference between the leading order and next-to-leading order matrix elements in the non-relativistic limit, summed and averaged over the final and initial polarizations,
@@
We compute the one loop diagrams in the non-relativistic limit with the aid of FeynCalc Shtabovenko et al. (2020), interfere them with the LO matrix elements, and reduce the resulting expressions via the Passarino-Veltman procedure Passarino and Veltman (1979) to scalar integrals which we evaluate numerically using LoopTools Hahn and Perez-Victoria (1999). We cross-checked our results by hand as well as by using Package-X Patel (2017). We find that the ultraviolet divergences cancel between the DM wave function corrections (see appendix A) and the vertex correction, as is expected based on the analogue of the Ward identity for U(1). UV divergences in the one-loop correction to the dark photon propagator cancel against the $\delta Z_{A^{\prime}}$ and $\delta m_{A^{\prime}}$ counter-terms, renormalized in the on-shell scheme (see appendix B for details), resulting in a final expression that is finite.
The resulting correction to the annihilation cross section is expressed as
$\langle\delta\sigma v\rangle_{\text{NLO}}$ $=-\frac{\alpha_{D}\,\langle\sigma v\rangle_{\text{LO}}}{3\,\pi}-\frac{\alpha_{D}\,\langle\sigma v\rangle_{\text{LO}}}{6\,\pi\,m_{\chi}^{2}\left(4\,m_{\chi}^{2}-m_{A^{\prime}}^{2}\right)}\Bigg{\{}-2\left(m_{A^{\prime}}^{2}-4m_{\chi}^{2}\right)$ (10)
@@
$+\left(m_{A^{\prime}}^{4}+4\,m_{A^{\prime}}^{2}m_{\chi}^{2}-32\,m_{\chi}^{4}\right)B_{0}\left(4\,m_{\chi}^{2};m_{\chi}^{2},m_{\chi}^{2}\right)-24\,m_{\chi}^{4}\,B_{0}\left(4\,m_{\chi}^{2};m_{\chi}^{2},m_{\chi}^{2}\right)$
@@
where $A_{0}$, $B_{0}$, and $C_{0}$ are scalar integral functions, and $B^{\prime}_{0}$ is the derivative of $B_{0}$ with respect to its first argument. We find that the contribution of the dark scalar to the dark photon self energy is quantitatively quite small and thus does not change the results by much, whereas the analytic expression is rather complicated. Hence, for brevity we omit it from the expression above. The full annihilation cross section to ${\cal O}(\alpha_{D}^{2})$ is:
$\langle\sigma v\rangle=\langle\sigma v\rangle_{\text{LO}}+\langle\delta\sigma v\rangle_{\text{NLO}}+{\cal O}\left(\alpha_{D}^{3}\right).$ (11)
Figure 3: The ratio $\langle\delta\sigma v\rangle_{\text{NLO}}/(\alpha_{D}\langle\sigma v\rangle_{\text{LO}})$ as a function of $m_{A^{\prime}}/m_{\chi}$, for $m_{\chi}=1$ MeV and for different ratios of $m_{H}/m_{A^{\prime}}=0.1,\,0.25\text{ and }1$, represented by the blue, magenta and red lines, respectively.
The impact of the one-loop correction can be summarized by the quantity $\langle\delta\sigma v\rangle_{\text{NLO}}/(\alpha_{D}\langle\sigma v\rangle_{\text{LO}})$, which characterizes the relative change compared to the leading order cross section with the $\alpha_{D}$ dependence scaled out. In Figure 3, we plot this ratio as a function of the ratio of $m_{A^{\prime}}/m_{\chi}$, for $m_{\chi}=1$ MeV for different values of $m_{H}/m_{A^{\prime}}$. For $m_{H}\geq m_{A^{\prime}}$, the inclusion of the dark Higgs negligibly impacts the annihilation cross-section, whereas in the light dark Higgs regime, $m_{H}\ll m_{A^{\prime}}$, the contribution is positive, though subdominant compared to the contributions from virtual dark photons. These results indicate that for $m_{A^{\prime}}/m_{\chi}\lesssim 4$, the corrections are positive, and can be as large as tens of percent near resonance where $m_{A^{\prime}}\sim 2\leavevmode\nobreak\ m_{\chi}$. For larger mass ratios, the impact of the higher order corrections is to decrease the net annihilation rate by an amount that flattens out for very large mass ratios. We find that, provided $m_{A^{\prime}}$ and $m_{H}$ are specified as a ratio to $m_{\chi}$, the quantity $\langle\delta\sigma v\rangle_{\text{NLO}}/(\alpha_{D}\langle\sigma v\rangle_{\text{LO}})$ is very insensitive to the mass of the dark matter itself over the entire range of interest.

### v2

@@ Abstract
Light thermal dark matter has emerged as an attractive theoretical possibility and a promising target for discovery at experiments in the near future. Such scenarios generically invoke mediators with very small couplings to the Standard Model, but with moderately strong couplings within the dark sector, calling into question theoretical estimates based on the lowest order of perturbation theory. As an example, we focus on a scenario in which (pseudo-)Dirac fermion dark matter is connected to the Standard Model via a dark photon charged under a new $U(1)^{\prime}$ extension of the standard model, and we investigate the impact of the next-to-leading order corrections to annihilation and scattering. We find that radiative corrections can significantly impact model predictions for the relic density and scattering cross-section, depending on the strength of the dark sector coupling and ratio of the dark matter to mediator mass. We also show why factorization into the yield parameter $Y$ typically presented in literature leads to imprecision. Our results are necessary to accurately map experimental searches into the model parameter space and assess their ability to reach thermal production targets.
@@ I Introduction
Of all of the viable scenarios of dark matter (DM), one of the most compelling is that of particles that are in thermal equilibrium with the Standard Model (SM) bath in the early Universe. DM with a thermal history is very well motivated because it not only provides a feasible prediction for large non-gravitational interactions between dark and ordinary matter, but it is also highly predictive in nature, leading to clear targets for experimental searches. Generically, after the DM has frozen out and its relic density set, it maintains the same interaction strength with the SM today, predicting that it is likely to be observable at a variety of experiments on Earth Bertone and Tait (2018). A canonical example is that of weakly interacting massive particles (WIMPs), a class of heavy ($\sim$ GeV-TeV scale) particles that interact with the SM through a roughly electroweak strength force Cushman et al. (2013); Cooley et al. (2022).
@@
However, despite decades of search, heavy ($\gtrsim$ GeV) WIMPs are yet to be discovered and their parameter space has become tightly constrained Cooley et al. (2022). This lack of a clear signal has motivated searches for other visions of DM, including models with masses in the MeV to GeV regime interacting with the SM via new undiscovered forces. While this class of DM is no longer a WIMP in the traditional sense, it can still be in equilibrium at early times and thus represents a thermal target. It is a very attractive prospect as it may provide a road map to a wide and rich dark sector. The possibilities of simplified DM models are vast. However, one of the most appealing scenarios is that of DM interacting with the SM through a new vector boson, often referred to as a dark photon, corresponding to a $U(1)^{\prime}$ extension of the SM Fayet (1980, 1981); Holdom (1986).
@@
Thermal sub-GeV DM, through a dark photon portal is very well motivated and has been the subject of much exploration in the literature. In this scenario, MeV to GeV mass DM can be produced and detected at a variety of current and near-future low energy accelerator experiments Izaguirre et al. (2015a, b, c, 2016); Feng and Smolinsky (2017); Izaguirre et al. (2017); Aguilar-Arevalo et al. (2018); Marsicano et al. (2018); Berlin and Kling (2019); Mohlabeng (2019); Tsai et al. (2021); Duerr et al. (2020); Berlin et al. (2020); Batell et al. (2021a, b); Carrillo González and Toro (2022); Cazzaniga et al. (2021); Gori et al. (2022); Krnjaic et al. (2022); Mongillo et al. (2023); Abdullahi et al. (2023); Brahma et al. (2023). Moreover, ambient DM in the Solar System can scatter with electrons and nucleons in small scale direct detection experiments with low energy recoil thresholds Emken et al. (2019); Bloch et al. (2021); Essig et al. (2022). In addition, accelerator and underground detector probes are well complemented by astrophysical and cosmological constraints An et al. (2018); Chang et al. (2018); DeRocco et al. (2019); Fiorillo and Vitagliano (2024); Fitzpatrick et al. (2022); Giovanetti et al. (2022). All together, these probes make for an exciting experimental program that shows promise for discovering light thermal DM in the near future.
@@
As an illustrative model, we focus on pseudo-Dirac dark matter interacting with the SM via a dark photon mediator. While we present results for the case in which the dark photon’s interactions with the SM arise entirely via kinetic mixing with the hypercharge boson, our results are actually more general and apply to any theory in which a pseudo-Dirac dark matter particle interacts with a vector mediator whose interactions with the SM are much weaker than those with the dark matter itself. Current experimental constraints and projections on this model that target the dark matter thermal relic abundance focus on a dark sector coupling $\alpha_{D}$ that is stronger than the QCD coupling at the electroweak scale. As a result, the dark sector is strongly coupled and relatively large corrections from higher orders of perturbation theory are expected. In this work, we compute the next-to-leading (NLO) order corrections on both the thermal annihilation and late time scattering processes with electrons. We focus on the NLO corrections at ${\cal O}(\alpha_{D})$ and neglect the presumably negligible higher order corrections from the tiny kinetic mixing.
@@
We find that the NLO corrections can be as large as ${\cal O}(10\%)$ for parameters typically discussed in the literature, and are thus necessary to take into account when precisely mapping experimental searches into the parameter space, and when comparing them with thermal production milestones that serve as a prime target for a currently growing intensity frontier program. It is worth pointing out that various theoretical constraints on this model resulting from its strong coupling have been considered in the literature, e.g. from the running of the dark sector coupling and the breakdown of perturbation theory Davoudiasl and Marciano (2015); Reilly and Toro (2023).
@@
The rest of this article is organized as follows: In Sec. II we describe the reference model, Sec. III is devoted to a full description of the thermal annihilation cross-section, including both leading and next-to-leading order (NLO) processes. We compute the scattering cross-sections with electrons relevant for direct detection searches in Sec. IV. Finally we conclude in Sec. V. The appendices provide some technical details related to the counterterms in the on-shell renormalization scheme.
@@ II Pseudo-Dirac Dark Matter
The basic module we consider consists of two Weyl fermions which play the role of dark matter and are paired by a Dirac mass, $m_{\chi}$. These fermions are neutral under the SM gauge groups but have equal and opposite charge ±$\pm 1$ under a gauged dark U(1), with corresponding gauge boson $A^{\prime}_{\mu}$. The U(1) symmetry is spontaneously broken by the vacuum expectation value of a dark scalar $\phi$, generating a mass for $A^{\prime}_{\mu}$ and [assuming the $\phi$ charge under U(1) is chosen appropriately] Majorana masses for the two Weyl fermions. We further assume that there is kinetic mixing between the dark photon and the SM hypercharge interaction, induced by unspecified UV physics.
@@
where $\chi$ is a Dirac fermion packaging both of the original Weyl fermions in the limit of zero mass splitting (and thus including both the dark matter and its heavier partner when the mass splitting is taken into account). $D_{\mu}\equiv\partial_{\mu}-ig_{D}q_{A^{\prime}}A^{\prime}_{\mu}$ is the covariant derivative for field of U(1) charge $q_{A^{\prime}}$, $X_{\mu\nu}$ and $B_{\mu\nu}$ are the field strengths for the U(1) and SM hypercharge bosons, respectively, and $\varepsilon$ characterizes the strength of the kinetic mixing. We normalize $q_{A^{\prime}}=1$ from here on for $\chi$. After symmetry breaking, the scalar field acquires a vacuum expectation value (VEV) $v_{\phi}$ and can be parametrized in the unitary gauge as $\phi\rightarrow(v_{\phi}+H_{D})/\sqrt{2}$, where $H_{D}$ is the dark Higgs boson.
@@ III Thermal Freeze-out
For dark matter masses below $\lesssim$ GeV, the annihilation rate into hadrons may be inferred by making use of the fact that $A^{\prime}$ couples to the same electromagnetic current as the photon. Thus, the effective cross section for annihilation into hadrons at a given center-of-mass energy can be related to the ratio $R$ measured in $e^{+}e^{-}$ reactions Olive et al. (2014):
@@
III.1 Tree-level annihilation
At tree level, the annihilation into SM fermions proceeds via $s$-channel exchange of the $A^{\prime}$ (see Fig. 1). The leading order matrix element $\overline{|{\cal M}|^{2}}_{\text{LO}}$ is given by
@@ III.1 Tree-level annihilation
$\Gamma(A^{\prime}\rightarrow\chi\bar{\chi})=\frac{\alpha_{D}}{3}\,m_{A^{\prime}}\left(1+\frac{2m_{\chi}^{2}}{m_{A^{\prime}}^{2}}\right)\,\sqrt{1-\frac{4m_{\chi}^{2}}{m_{A^{\prime}}^{2}}},$ (6)
$\Gamma(A^{\prime}\rightarrow f\bar{f})=\frac{\alpha Q_{f}^{2}}{3}\,\varepsilon^{2}\,m_{A^{\prime}}\left(1+\frac{2m_{f}^{2}}{m_{A^{\prime}}^{2}}\right)\,\sqrt{1-\frac{4m_{f}^{2}}{m_{A^{\prime}}^{2}}},$ (7)
@@
In the non-relativistic limit, the thermally averaged tree-level cross section is
@@ III.2 NLO corrections
At ${\cal O}(\alpha_{D}^{2})$, the annihilation cross section receives virtual corrections in the form of self-energy corrections to the dark photon propagator (these include diagrams with fermions and the dark Higgs in the loop) and incoming dark matter wave functions as well as a correction to the $A^{\prime}$-$\chi$-$\overline{\chi}$ vertex (see Fig. 2). In the regime of $m_{A}^{\prime}\geq 2m_{\chi}$, an additional emitted $A^{\prime}$ would always be off-shell, leading the real emission corrections to be effectively higher order in $\varepsilon$ such that they can be safely neglected. The ${\cal O}(\alpha_{D}^{2})$ correction to the annihilation process is given by the interference between the leading order and next-to-leading order matrix elements in the non-relativistic limit, summed and averaged over the final and initial polarizations,
@@
We compute the one loop diagrams in the non-relativistic limit with the aid of FeynCalc Shtabovenko et al. (2020), interfere them with the LO matrix elements, and reduce the resulting expressions via the Passarino-Veltman procedure Passarino and Veltman (1979) to scalar integrals which we evaluate numerically using LoopTools Hahn and Perez-Victoria (1999). We cross-checked our results by hand as well as by using Package-X Patel (2017). We find that the ultraviolet divergences cancel between the DM wave function corrections (see Appendix A) and the vertex correction, as is expected based on the analogue of the Ward identity for U(1). UV divergences in the one-loop correction to the dark photon propagator cancel against the $\delta Z_{A^{\prime}}$ and $\delta m_{A^{\prime}}^{2}$ counterterms, renormalized in the on-shell scheme (see Appendix B for details), resulting in a final expression that is finite.
The resulting correction to the annihilation cross section from diagrams without a dark Higgs is expressed as
$\langle\delta\sigma v\rangle_{\text{NLO}}^{\text{DM}}$ $=-\frac{\alpha_{D}\,\langle\sigma v\rangle_{\text{LO}}}{3\,\pi}-\frac{\alpha_{D}\,\langle\sigma v\rangle_{\text{LO}}}{6\,\pi\,m_{\chi}^{2}\left(4\,m_{\chi}^{2}-m_{A^{\prime}}^{2}\right)}\Bigg{\{}-2\left(m_{A^{\prime}}^{2}-4m_{\chi}^{2}\right)$ (10)
@@
$+\left(m_{A^{\prime}}^{4}+4\,m_{A^{\prime}}^{2}m_{\chi}^{2}-56\,m_{\chi}^{4}\right)B_{0}\left(4\,m_{\chi}^{2};m_{\chi}^{2},m_{\chi}^{2}\right)$
@@
where $A_{0}$, $B_{0}$, and $C_{0}$ are scalar integral functions, and $B^{\prime}_{0}$ is the derivative of $B_{0}$ with respect to its first argument. The contribution from the one-loop correction involving the dark Higgs is
$\langle\delta\sigma v\rangle_{\text{NLO}}^{\text{Higgs}}$ $=\frac{\langle\sigma v\rangle_{LO}\,\alpha_{D}\,q_{\phi}^{2}}{24\,\pi\,m_{A^{\prime}}^{4}m_{\chi}^{2}\left(m_{A^{\prime}}^{2}-4m_{\chi}^{2}\right)}\Bigg{\{}B_{0}\left(m_{A^{\prime}}^{2};m_{A^{\prime}}^{2},m_{H}^{2}\right)\Bigg{[}-8m_{\chi}^{2}\Bigg{(}6m_{A^{\prime}}^{6}$ (11)
$-3m_{A^{\prime}}^{4}m_{H}^{2}+m_{A^{\prime}}^{2}\left(m_{H}^{4}+4m_{H}^{2}m_{\chi}^{2}\right)-2m_{H}^{4}m_{\chi}^{2}\Bigg{)}\Bigg{]}$
$+B_{0}\left(4m_{\chi}^{2};m_{A^{\prime}}^{2},m_{H}^{2}\right)\Bigg{[}m_{A^{\prime}}^{4}\left(m_{A^{\prime}}^{4}-2m_{A^{\prime}}^{2}\left(m_{H}^{2}-20m_{\chi}^{2}\right)+\left(m_{H}^{2}-4m_{\chi}^{2}\right)^{2}\right)\Bigg{]}$
$+B_{0}^{\prime}\left(m_{A^{\prime}}^{2};m_{A^{\prime}}^{2},m_{H}^{2}\right)\Bigg{[}4m_{A^{\prime}}^{2}m_{\chi}^{2}\left(12m_{A^{\prime}}^{4}-4m_{A^{\prime}}^{2}m_{H}^{2}+m_{H}^{4}\right)\left(m_{A^{\prime}}^{2}-4m_{\chi}^{2}\right)\Bigg{]}$
$+(m_{A^{\prime}}^{2}-m_{H}^{2})\left(m_{A^{\prime}}^{2}-4m_{\chi}^{2}\right)^{2}\Bigg{[}A_{0}\left(m_{H}^{2}\right)-A_{0}\left(m_{A^{\prime}}^{2}\right)\Bigg{]}\Bigg{\}}$
Figure 3: The ratio $\langle\delta\sigma v\rangle_{\text{NLO}}/(\alpha_{D}\langle\sigma v\rangle_{\text{LO}})$ as a function of $m_{A^{\prime}}/m_{\chi}$, for $m_{\chi}=1$ MeV and for different ratios of $m_{H}/m_{A^{\prime}}=0.1,\,0.25,1\text{ and }10$, represented by the blue, magenta, red and orange lines, respectively.
The full annihilation cross section to ${\cal O}(\alpha_{D}^{2})$ is
$\langle\sigma v\rangle=\langle\sigma v\rangle_{\text{LO}}+\langle\delta\sigma v\rangle_{\text{NLO}}^{\text{DM}}+\langle\delta\sigma v\rangle_{\text{NLO}}^{\text{Higgs}}+{\cal O}\left(\alpha_{D}^{3}\right).$ (12)
The impact of the one-loop correction can be summarized by the quantity $\langle\delta\sigma v\rangle_{\text{NLO}}/(\alpha_{D}\langle\sigma v\rangle_{\text{LO}})$, which characterizes the relative change compared to the leading order cross section with the $\alpha_{D}$ dependence scaled out. In Fig. 3, we plot this ratio as a function of the ratio of $m_{A^{\prime}}/m_{\chi}$, for $m_{\chi}=1$ MeV for different values of $m_{H}/m_{A^{\prime}}$. As illustrated in the figure, in the limit of large dark Higgs mass, the net effect of the NLO corrections is to reduce the cross section by a modest amount. As $m_{H}/m_{A^{\prime}}$ decreases, the NLO effect is to increase the total annihilation cross-section significantly. This trend saturates around $m_{A^{\prime}}/m_{\chi}\sim 5$, after which the total cross-section flattens out for very large mass ratios. We also find that, provided $m_{A^{\prime}}$ and $m_{H}$ are specified as a ratio to $m_{\chi}$, the quantity $\langle\delta\sigma v\rangle_{\text{NLO}}/(\alpha_{D}\langle\sigma v\rangle_{\text{LO}})$ is insensitive to the mass of the dark matter over the entire range of interest, while we choose a benchmark $m_{\chi}=1$ MeV, the results are identical for e.g., $m_{\chi}=1$ GeV.